# Consensus over Clustered Networks Using Intermittent and Asynchronous Output Feedback

### v2

@@ V Stability Analysis
$T_{\min}$ $\triangleq\min\{T_{1}^{p}\colon p\in\mathcal{V}\}\cup\{T_{3}^{r}\colon r\in[M^{\ast}]\},$
$T_{\max}$ $\triangleq\max\{T_{2}^{p}\colon p\in\mathcal{V}\}\cup\{T_{4}^{r}\colon r\in[M^{\ast}]\}.$
@@
    $\left(\frac{j}{N+M^{\ast}}-1\right)T_{\min}\leq t\leq\frac{j}{N+M^{\ast}}T_{\max}.$ (29)
@@
$T_{2}$ $\triangleq(T_{2}^{p})_{p\in\mathcal{V}}\in\mathbb{R}^{N},\quad T_{4}\triangleq(T_{4}^{r})_{r\in[M^{\ast}]}\in\mathbb{R}^{M^{\ast}},$
$T_{2}^{\max}$ $\triangleq\max\{T_{2}^{p}\colon p\in\mathcal{V}\}\in\mathbb{R}_{>0}.$
@@
$P_{2}(\tau)$ $\triangleq\text{diag}(P_{2,1}\exp(\sigma\tau_{1}),P_{2,2}\exp(\sigma\tau_{2}),...,$ (30)
$\phantom{\triangleq}\ P_{2,N}\exp(\sigma\tau_{N}))\in\mathbb{R}^{mN\times mN},$
$P_{3}(\rho)$ $\triangleq\text{diag}(P_{3,1}\exp(\sigma\rho_{1}),P_{3,2}\exp(\sigma\rho_{2}),...,$
$\phantom{\triangleq}\ P_{3,M^{\ast}}\exp(\sigma\rho_{M^{\ast}}))\in\mathbb{R}^{mNM^{\ast}\times mNM^{\ast}},$
$P(\tau,\rho)$ $\triangleq\text{diag}(P_{1},P_{2}(\tau),P_{3}(\rho)).$
@@
$\mathsf{M}(\tau,\rho)$ $\triangleq\mathsf{F}^{\top}P(\tau,\rho)+P(\tau,\rho)\mathsf{F}+\dot{P}(\tau,\rho),$ (31)
$\mu$ $\triangleq-\sup\{\lambda_{\max}(\mathsf{M}(\tau,\rho))\colon(\tau,\rho)\in\mathcal{T}\times\mathcal{R}\},$
$\alpha_{1}$ $\triangleq\lambda_{\min}(P(0_{N},0_{M^{\ast}})),\ \alpha_{2}\triangleq\lambda_{\max}(P(T_{2},T_{4})),$
$\dot{P}(\tau,\rho)$ $=-\sigma\text{diag}(0_{n(N-1)\times n(N-1)},P_{2}(\tau),P_{3}(\rho)).$
By definition, the matrix $\mathsf{M}$ is symmetric. If $\mathsf{M}$ satisfies $\mathsf{M}<\mathbf{0}$ for all $(\tau,\rho)\in\mathcal{T}\times\mathcal{R}$, then the constant $\mu$ is positive. Further, $\alpha_{1}\in\mathbb{R}_{>0}$ and $\alpha_{2}\in\mathbb{R}_{>0}$ since $P(\tau,\rho)$ is positive definite.
@@
$\kappa$ $\triangleq\sqrt{\frac{\alpha_{2}}{\alpha_{1}}\exp\left(\frac{(1-\varepsilon)\mu T_{\min}}{\alpha_{2}}\right)}\in\mathbb{R}_{>0},$
$\alpha$ $\triangleq\frac{1}{2}\min\left\{\frac{\varepsilon\mu}{\alpha_{2}},\frac{(1-\varepsilon)\mu T_{\min}}{\alpha_{2}(N+M^{\ast})}\right\}\in\mathbb{R}_{>0},$
@@
$\dot{V}(\xi)$ $=2z^{\top}P(\tau,\rho)\mathsf{F}z+z^{\top}\dot{P}(\tau,\rho)z$ (35)
$=z^{\top}\big{(}\mathsf{F}^{\top}P(\tau,\rho)+P(\tau,\rho)\mathsf{F}+\dot{P}(\tau,\rho)\big{)}z$
$=z^{\top}\mathsf{M}(\tau,\rho)z.$
@@
When $\xi\in D$ and $g\in G(\xi)$, the change in $V(\xi)$ is computed using $\Delta V(\xi)=V(g)-V(\xi)$. Suppose $\tau_{q}=0$ and $\rho_{s}=0$ for some $q\in\mathcal{V}$ and $s\in[M^{\ast}]$ without loss of generality. Hence, the jump map in (26) implies that [^7] The result in (37) also follows if $\tau_{q}=0$, $\rho_{s}=0$, or any combination of timers take on a value of 00.
$=-\tilde{\eta}_{q}^{\top}P_{2,q}\tilde{\eta}_{q}-\tilde{\zeta}_{s}^{\top}P_{3,s}\tilde{\zeta}_{s}\leq 0$
@@
dom $\text{dom }\phi\bigcap([0,t_{j}]\times\{0,1,...,j-1\})=\bigcup_{k=1}^{j}([t_{k-1},t_{k}]\times\{k-1\}).$
@@
$-\frac{\mu}{\alpha_{2}}t\leq-\min\left\{\frac{\varepsilon\mu}{\alpha_{2}},\frac{(1-\varepsilon)\mu T_{\min}}{\alpha_{2}(N+M^{\ast})}\right\}(t+j)+\frac{(1-\varepsilon)\mu T_{\min}}{\alpha_{2}},$ (42)
@@
$V(\phi(t,j))$ $\leq\exp\left(-\min\left\{\frac{\varepsilon\mu}{\alpha_{2}},\frac{(1-\varepsilon)\mu T_{\min}}{\alpha_{2}(N+M^{\ast})}\right\}(t+j)\right)$ (43)
$\phantom{=}\ \cdot\exp\left(\frac{\left(1-\varepsilon\right)\mu T_{\min}}{\alpha_{2}}\right)V\left(\phi\left(0,0\right)\right).$
@@
$|\phi(t,j)|_{\mathcal{A}}^{2}$ $\leq\frac{\alpha_{2}}{\alpha_{1}}\exp\left(-\min\left\{\frac{\varepsilon\mu}{\alpha_{2}},\frac{(1-\varepsilon)\mu T_{\min}}{\alpha_{2}(N+M^{\ast})}\right\}(t+j)\right)$ (44)
$\phantom{=}\ \cdot\exp\left(\frac{\left(1-\varepsilon\right)\mu T_{\min}}{\alpha_{2}}\right)|\phi(0,0)|_{\mathcal{A}}^{2},$
@@
Using (33) and (43), one can conclude that $z=z(\phi(t,j))$ is bounded for all $(t,j)\in\text{dom }\phi$. Recall $(\mathsf{V}^{\top}\otimes I_{n})$ is full rank. Thus, $z=(x^{\circ},\tilde{\eta},\tilde{\zeta})$ implies $x_{p}$, $\tilde{\eta}_{p}$, and $\tilde{\zeta}_{pr}$ are bounded for all $p\in\mathcal{V}$ and all $r\in[M^{\ast}]$. Using the output equation in (5), $x_{p}$ being bounded implies $y_{p}$ is bounded for all $p\in\mathcal{V}$. Hence, $u_{p}$ is bounded for all $p\in\mathcal{V}$ given (8)–(10) and the bounded flow intervals of $\mathcal{H}$, i.e., $t_{j+1}-t_{j}\leq T_{\max}$ for all $(t_{j})_{j=0}^{\infty}$. ∎
With respect to Theorem 1, the global exponential stability of $\mathcal{A}$ for $\mathcal{H}$ requires $\mathsf{M}(\tau,\rho)<\mathbf{0}$ to hold for all $(\tau,\rho)\in\mathcal{T}\times\mathcal{R}$. In other words, $\mathsf{M}(\tau,\rho)<\mathbf{0}$ must hold for an infinite number of points, which cannot be exhaustively verified in practice. Hence, the following result is motivated.
For each $p\in\mathcal{V}$ and $r\in[M^{\ast}]$, let $\gamma_{p}\colon[0,T_{2}^{p}]\to[0,1]$ and $\gamma^{r}\colon[0,T_{4}^{r}]\to[0,1]$, such that
$\gamma_{p}(w)$ $\triangleq\frac{\exp(\sigma w)-\exp(\sigma T_{2}^{p})}{1-\exp(\sigma T_{2}^{p})},$ (45)
$\gamma^{r}(w)$ $\triangleq\frac{\exp(\sigma w)-\exp(\sigma T_{4}^{r})}{1-\exp(\sigma T_{4}^{r})}.$
Also, let $\epsilon\in[0,1]$ and $\theta\triangleq(\epsilon,\tau,\rho)\in\Theta$, where $\Theta\triangleq[0,1]\times\mathcal{T}\times\mathcal{R}$. We can then construct the following block diagonal matrices:
$R_{2}(\tau)$ $\triangleq\text{diag}(\gamma_{1}(\tau_{1}),\gamma_{2}(\tau_{2}),...,\gamma_{N}(\tau_{N}))\otimes I_{m},$ (46)
$R_{3}(\rho)$ $\triangleq\text{diag}(\gamma^{1}(\rho_{1}),\gamma^{2}(\rho_{2}),...,\gamma^{M^{\ast}}(\rho_{M^{\ast}}))\otimes I_{mN},$
$R(\theta)$ $\triangleq\text{diag}(\epsilon I_{n(N-1)},R_{2}(\tau),R_{3}(\rho)),$
where $R_{2}(\tau)\in\mathbb{R}^{mN\times mN}$ and $R_{3}(\rho)\in\mathbb{R}^{mNM^{\ast}\times mNM^{\ast}}$.
@@
If there exist timer parameters $0<T_{1}^{p}\leq T_{2}^{p}$ for each agent $p\in\mathcal{V}$ and $0<T_{3}^{r}\leq T_{4}^{r}$ for each $r\in[M^{\ast}]$, a constant $\sigma$, controller matrix $K_{u}$, estimator matrices $K_{\eta,p}$ for each $p\in\mathcal{V}$, estimator matrices $K_{\zeta,pr}$ for every $(p,r)\in\mathcal{V}\times[M^{\ast}]$, and symmetric, positive definite matrices $P_{1}$, $P_{2,p}$ for every $p\in\mathcal{V}$, and $P_{3,r}$ for every $r\in[M^{\ast}]$ that satisfy $\mathsf{M}(0_{N},0_{M^{\ast}})<\mathbf{0}$ and $\mathsf{M}(T_{2},T_{4})<\mathbf{0}$, then $\mathsf{M}(\tau,\rho)<\mathbf{0}$ for all $(\tau,\rho)\in\mathcal{T}\times\mathcal{R}$. $\triangle$
@@
For all $p\in\mathcal{V}$ and $r\in[M^{\ast}]$, the functions in (45) can be written as
$\exp(\sigma\tau_{p})$ $=\gamma_{p}(\tau_{p})+(1-\gamma_{p}(\tau_{p}))\exp(\sigma T_{2}^{p}),$ (47)
$\exp(\sigma\rho_{r})$ $=\gamma^{r}(\rho_{r})+(1-\gamma^{r}(\rho_{r}))\exp(\sigma T_{4}^{r}).$
Using the equations in (47) and the definitions of $P(\tau,\rho)$ and $R(\theta)$, one can derive
$P(\tau,\rho)$ $=R(\theta)P(0_{N},0_{M^{\ast}})+(I_{\varrho}-R(\theta))P(T_{2},T_{4}),$ (48)
$\dot{P}(\tau,\rho)$ $=R(\theta)\dot{P}(0_{N},0_{M^{\ast}})+(I_{\varrho}-R(\theta))\dot{P}(T_{2},T_{4}),$
where $\varrho\triangleq n(N-1)+mN+mNM^{\ast}$. Let $v$ be an arbitrary real-valued vector such that the quadratic form $v^{\top}\mathsf{M}(\tau,\rho)v$ is well-defined. Using the definition of $\mathsf{M}$ in (31), one can obtain
$v^{\top}\mathsf{M}(\tau,\rho)v=2v^{\top}P(\tau,\rho)\mathsf{F}v+v^{\top}\dot{P}(\tau,\rho)v.$ (49)
@@
$v^{\top}\mathsf{M}(\tau,\rho)v$ $=v^{\top}R(\theta)(2P(0_{N},0_{M^{\ast}})\mathsf{F}+\dot{P}(0_{N},0_{M^{\ast}}))v$ (50)
$\phantom{=}\ +v^{\top}(I_{\varrho}-R(\theta))(2P(T_{2},T_{4})\mathsf{F}+\dot{P}(T_{2},T_{4}))v.$
Since the matrices $P(\tau,\rho)$, $\dot{P}(\tau,\rho)$, and $R(\theta)$ are symmetric, (50) can be simplified to
$v^{\top}\mathsf{M}(\tau,\rho)v$ $=v^{\top}R(\theta)\mathsf{M}(0_{N},0_{M^{\ast}})R(\theta)v$ (51)
$\phantom{=}\ +v^{\top}(I_{\varrho}-R(\theta))\mathsf{M}(T_{2},T_{4})(I_{\varrho}-R(\theta))v.$
Although the matrices $R(\theta)$ and $I_{\varrho}-R(\theta)$ are positive semi-definite for all $\theta\in\Theta$, $R(\theta)v$ and $(I_{\varrho}-R(\theta))v$ cannot be zero simultaneously for any $\theta\in\Theta$ and $v\neq\mathbf{0}$. Thus, for all $v\neq\mathbf{0}$, $\mathsf{M}(0_{N},0_{M^{\ast}})<\mathbf{0}$, $\mathsf{M}(T_{2},T_{4})<\mathbf{0}$, and the equality in (51) imply $\mathsf{M}(\tau,\rho)<\mathbf{0}$ for all $(\tau,\rho)\in\mathcal{T}\times\mathcal{R}$. ∎
Corollary 1 provides a convenient method for validating the sufficient condition of Theorem 1. Specifically, Corollary 1 only requires the checking of two matrix inequalities, namely, $\mathsf{M}(0_{N},0_{M^{\ast}})<\mathbf{0}$ and $\mathsf{M}(T_{2},T_{4})<\mathbf{0}$, to ensure $\mathsf{M}(\tau,\rho)<\mathbf{0}$ holds for all $(\tau,\rho)\in\mathcal{T}\times\mathcal{R}$.
VI Robustness
In this section, we briefly discuss the robustness properties of the set $\mathcal{A}$ for the hybrid system $\mathcal{H}$. Since $\mathcal{H}$ is nominally well-posed and the set $\mathcal{A}$ is compact and GES for $\mathcal{H}$, standard arguments can be used to show that $\mathcal{A}$ is robust to vanishing perturbations for $\mathcal{H}$ (for example, see [32, Section 7.3] and [38, Section 3.5.1]). The reader is referred to [32, Chapter 7] for results showcasing the robustness of $\mathcal{A}$ for $\mathcal{H}$, which is due to the structure of $\mathcal{A}$ and $\mathcal{H}$ and the exponential stability of $\mathcal{A}$. For the remainder of this section, we examine the robustness of $\mathcal{A}$ to non-vanishing measurement noise injected via a modified jump map of $\mathcal{H}$.

### v3

@@ 5 Stability Analysis
$T_{\min}\coloneqq\min\{T_{1}^{p}\colon p\in\mathcal{V}\}\cup\{T_{3}^{r}\colon r\in[M^{\ast}]\},\quad T_{\max}\coloneqq\max\{T_{2}^{p}\colon p\in\mathcal{V}\}\cup\{T_{4}^{r}\colon r\in[M^{\ast}]\}.$
@@
    $\left(\frac{j}{N+M^{\ast}}-1\right)T_{\min}\leq t\leq\frac{j}{N+M^{\ast}}T_{\max}$ (30)
@@
$T_{2}\coloneqq(T_{2}^{p})_{p\in\mathcal{V}}\in\mathbb{R}^{N},\quad T_{4}\coloneqq(T_{4}^{r})_{r\in[M^{\ast}]}\in\mathbb{R}^{M^{\ast}},\quad T_{2}^{\max}\coloneqq\max\{T_{2}^{p}\colon p\in\mathcal{V}\}\in\mathbb{R}_{>0}.$
@@
$P_{2}(\tau)$ $\coloneqq\text{diag}(P_{2,1}\exp(\sigma\tau_{1}),P_{2,2}\exp(\sigma\tau_{2}),.,P_{2,N}\exp(\sigma\tau_{N}))\in\mathbb{R}^{mN\times mN},$ (31)
$P_{3}(\rho)$ $\coloneqq\text{diag}(P_{3,1}\exp(\sigma\rho_{1}),P_{3,2}\exp(\sigma\rho_{2}),.,P_{3,M^{\ast}}\exp(\sigma\rho_{M^{\ast}}))\in\mathbb{R}^{mNM^{\ast}\times mNM^{\ast}},$
$P(\tau,\rho)$ $\coloneqq\text{diag}(P_{1},P_{2}(\tau),P_{3}(\rho)).$
@@
$\mathsf{M}(\tau,\rho)$ $\coloneqq\mathsf{F}^{\top}P(\tau,\rho)+P(\tau,\rho)\mathsf{F}+\dot{P}(\tau,\rho),\quad\mu\coloneqq-\sup\{\lambda_{\max}(\mathsf{M}(\tau,\rho))\colon(\tau,\rho)\in\mathcal{T}\times\mathcal{R}\},$ (32)
$\alpha_{1}$ $\coloneqq\lambda_{\min}(P(0_{N},0_{M^{\ast}})),\quad\alpha_{2}\coloneqq\lambda_{\max}(P(T_{2},T_{4})),\quad\dot{P}(\tau,\rho)=-\sigma\text{diag}(0_{n(N-1)\times n(N-1)},P_{2}(\tau),P_{3}(\rho)).$
By definition, the matrix $\mathsf{M}(\tau,\rho)$ is symmetric. If $\mathsf{M}(\tau,\rho)$ satisfies $\mathsf{M}(\tau,\rho)<\mathbf{0}$ for all $(\tau,\rho)\in\mathcal{T}\times\mathcal{R}$, then the constant $\mu$ is positive. Further, $\alpha_{1}\in\mathbb{R}_{>0}$ and $\alpha_{2}\in\mathbb{R}_{>0}$ since $P(\tau,\rho)$ is positive definite.
@@
$\kappa\coloneqq\sqrt{\frac{\alpha_{2}}{\alpha_{1}}\exp\left(\frac{(1-\varepsilon)\mu T_{\min}}{\alpha_{2}}\right)}\in\mathbb{R}_{>0},\quad\alpha\coloneqq\frac{1}{2}\min\left\{\frac{\varepsilon\mu}{\alpha_{2}},\frac{(1-\varepsilon)\mu T_{\min}}{\alpha_{2}(N+M^{\ast})}\right\}\in\mathbb{R}_{>0},$
@@
$\dot{V}(\xi)=2z^{\top}P(\tau,\rho)\mathsf{F}z+z^{\top}\dot{P}(\tau,\rho)z=z^{\top}\big(\mathsf{F}^{\top}P(\tau,\rho)+P(\tau,\rho)\mathsf{F}+\dot{P}(\tau,\rho)\big)z=z^{\top}\mathsf{M}(\tau,\rho)z.$ (36)
@@
When $\xi\in D$ and $g\in G(\xi)$, the change in $V(\xi)$ is computed using $\Delta V(\xi)=V(g)-V(\xi)$. Suppose $\tau_{q}=0$ and $\rho_{s}=0$ for some $q\in\mathcal{V}$ and $s\in[M^{\ast}]$ without loss of generality. Hence, the jump map in (27) implies that [^10] The result in (38) also follows if $\tau_{q}=0$, $\rho_{s}=0$, or any combination of timers take on a value of 0.
$\Delta V(\xi)=(z^{+})^{\top}P(\tau^{+},\rho^{+})(z^{+})-z^{\top}P(\tau,\rho)z=-\tilde{\eta}_{q}^{\top}P_{2,q}\tilde{\eta}_{q}-\tilde{\zeta}_{s}^{\top}P_{3,s}\tilde{\zeta}_{s}\leq 0$ (38)
@@
$\text{dom }\phi\bigcap([0,t_{j}]\times\{0,1,...,j-1\})=\bigcup_{k=1}^{j}([t_{k-1},t_{k}]\times\{k-1\}).$
@@
$\frac{\mu}{\alpha_{2}}t\leq-\min\left\{\frac{\varepsilon\mu}{\alpha_{2}},\frac{(1-\varepsilon)\mu T_{\min}}{\alpha_{2}(N+M^{\ast})}\right\}(t+j)+\frac{(1-\varepsilon)\mu T_{\min}}{\alpha_{2}},$ (43)
@@
$V(\phi(t,j))\leq\exp\left(-\min\left\{\frac{\varepsilon\mu}{\alpha_{2}},\frac{(1-\varepsilon)\mu T_{\min}}{\alpha_{2}(N+M^{\ast})}\right\}(t+j)\right)\exp\left(\frac{\left(1-\varepsilon\right)\mu T_{\min}}{\alpha_{2}}\right)V\left(\phi\left(0,0\right)\right).$ (44)
@@
$|\phi(t,j)|_{\mathcal{A}}^{2}\leq\frac{\alpha_{2}}{\alpha_{1}}\exp\left(-\min\left\{\frac{\varepsilon\mu}{\alpha_{2}},\frac{(1-\varepsilon)\mu T_{\min}}{\alpha_{2}(N+M^{\ast})}\right\}(t+j)\right)\exp\left(\frac{\left(1-\varepsilon\right)\mu T_{\min}}{\alpha_{2}}\right)|\phi(0,0)|_{\mathcal{A}}^{2},$ (45)
@@
[…]
With respect to Theorem 1, the GES of $\mathcal{A}$ for $\mathcal{H}$ requires $\mathsf{M}(\tau,\rho)<\mathbf{0}$ to hold for all $(\tau,\rho)\in\mathcal{T}\times\mathcal{R}$. In other words, $\mathsf{M}(\tau,\rho)<\mathbf{0}$ must hold for an infinite number of points, which cannot be exhaustively verified in practice. Hence, we require a practical means by which to verify $\mathsf{M}(\tau,\rho)<\mathbf{0}$ holds for all $(\tau,\rho)\in\mathcal{T}\times\mathcal{R}$, which motivates the following result.
@@
If there exist timer parameters $0<T_{1}^{p}\leq T_{2}^{p}$ for each $p\in\mathcal{V}$ and $0<T_{3}^{r}\leq T_{4}^{r}$ for each $r\in[M^{\ast}]$, a constant $\sigma$, controller matrix $K_{u}$, estimator matrices $K_{\eta,p}$ for each $p\in\mathcal{V}$, estimator matrices $K_{\zeta,pr}$ for every $(p,r)\in\mathcal{V}\times[M^{\ast}]$, and symmetric, positive definite matrices $P_{1}$, $P_{2,p}$ for every $p\in\mathcal{V}$, and $P_{3,r}$ for every $r\in[M^{\ast}]$ that satisfy $\mathsf{M}(0_{N},0_{M^{\ast}})<\mathbf{0}$ and $\mathsf{M}(T_{2},T_{4})<\mathbf{0}$, then $\mathsf{M}(\tau,\rho)<\mathbf{0}$ for all $(\tau,\rho)\in\mathcal{T}\times\mathcal{R}$. $\triangle$
@@
The following argument is inspired by the proof of Proposition 3.9 in [Li.Phillips.ea2018]. For each $p\in\mathcal{V}$ and $r\in[M^{\ast}]$, let $\gamma_{\tau,p}\colon[0,T_{2}^{p}]\to[0,1]$ and $\gamma_{\rho,r}\colon[0,T_{4}^{r}]\to[0,1]$, such that
$\gamma_{\tau,p}(w)\coloneqq\frac{\exp(\sigma w)-\exp(\sigma T_{2}^{p})}{1-\exp(\sigma T_{2}^{p})},\quad\gamma_{\rho,r}(w)\coloneqq\frac{\exp(\sigma w)-\exp(\sigma T_{4}^{r})}{1-\exp(\sigma T_{4}^{r})}.$ (46)
Let $\epsilon\in[0,1]$ and $\theta\coloneqq(\epsilon,\tau,\rho)\in\Theta$, where $\Theta\coloneqq[0,1]\times\mathcal{T}\times\mathcal{R}$. For each $p\in\mathcal{V}$ and $r\in[M^{\ast}]$, the functions in (46) can be manipulated to obtain
$\exp(\sigma\tau_{p})=\gamma_{\tau,p}(\tau_{p})+(1-\gamma_{\tau,p}(\tau_{p}))\exp(\sigma T_{2}^{p}),\quad\exp(\sigma\rho_{r})=\gamma_{\rho,r}(\rho_{r})+(1-\gamma_{\rho,r}(\rho_{r}))\exp(\sigma T_{4}^{r}).$ (47)
One can then construct the following block diagonal matrices:
$R_{2}(\tau)$ $\coloneqq\text{diag}(\gamma_{\tau,1}(\tau_{1}),\gamma_{\tau,2}(\tau_{2}),.,\gamma_{\tau,N}(\tau_{N}))\otimes I_{m},\quad R_{3}(\rho)\coloneqq\text{diag}(\gamma_{\rho,1}(\rho_{1}),\gamma_{\rho,2}(\rho_{2}),.,\gamma_{\rho,{M^{\ast}}}(\rho_{M^{\ast}}))\otimes I_{mN},$ (48)
$R(\theta)$ $\coloneqq\text{diag}((1-\epsilon)I_{n(N-1)},R_{2}(\tau),R_{3}(\rho)),$
where $R_{2}(\tau)\in\mathbb{R}^{mN\times mN}$ and $R_{3}(\rho)\in\mathbb{R}^{mNM^{\ast}\times mNM^{\ast}}$. Using the definition of $P(\tau,\rho)$ in (31), the equalities in (47), and the definition of $R(\theta)$ in (48), one can derive
$P(\tau,\rho)=R(\theta)P(0_{N},0_{M^{\ast}})+(I_{\varrho}-R(\theta))P(T_{2},T_{4}),\quad\dot{P}(\tau,\rho)=R(\theta)\dot{P}(0_{N},0_{M^{\ast}})+(I_{\varrho}-R(\theta))\dot{P}(T_{2},T_{4}),$ (49)
where $\varrho\coloneqq n(N-1)+mN+mNM^{\ast}$. Let $v$ be an arbitrary real-valued vector such that the quadratic form $v^{\top}\mathsf{M}(\tau,\rho)v$ is well-defined. Using the definition of $\mathsf{M}(\tau,\rho)$ in (32), it follows that
$v^{\top}\mathsf{M}(\tau,\rho)v=v^{\top}\left(2P(\tau,\rho)\mathsf{F}+\dot{P}(\tau,\rho)\right)v.$ (50)
@@
$v^{\top}\mathsf{M}(\tau,\rho)v=v^{\top}R(\theta)(2P(0_{N},0_{M^{\ast}})\mathsf{F}+\dot{P}(0_{N},0_{M^{\ast}}))v+v^{\top}(I_{\varrho}-R(\theta))(2P(T_{2},T_{4})\mathsf{F}+\dot{P}(T_{2},T_{4}))v.$ (51)
Since the matrices $P(\tau,\rho)$, $\dot{P}(\tau,\rho)$, $R(\theta)$, and $I_{\varrho}-R(\theta)$ are symmetric, (51) can be simplified to
$v^{\top}\mathsf{M}(\tau,\rho)v=v^{\top}B_{1}(\theta)\mathsf{M}(0_{N},0_{M^{\ast}})B_{1}(\theta)v+v^{\top}B_{2}(\theta)\mathsf{M}(T_{2},T_{4})B_{2}(\theta)v,$ (52)
where
$B_{1,2}(\tau)$ $\coloneqq\text{diag}\left(\sqrt{\gamma_{\tau,1}(\tau_{1})},\sqrt{\gamma_{\tau,2}(\tau_{2})},...,\sqrt{\gamma_{\tau,N}(\tau_{N})}\right)\otimes I_{m},$
$B_{1,3}(\rho)$ $\coloneqq\text{diag}\left(\sqrt{\gamma_{\rho,1}(\rho_{1})},\sqrt{\gamma_{\rho,2}(\rho_{2})},...,\sqrt{\gamma_{\rho,{M^{\ast}}}(\rho_{M^{\ast}})}\right)\otimes I_{mN},$
$B_{1}(\theta)$ $\coloneqq\text{diag}\left(\sqrt{1-\epsilon}I_{n(N-1)},B_{1,2}(\tau),B_{1,3}(\rho)\right)$
$B_{2,2}(\tau)$ $\coloneqq\text{diag}\left(\sqrt{1-\gamma_{\tau,1}(\tau_{1})},\sqrt{1-\gamma_{\tau,2}(\tau_{2})},...,\sqrt{1-\gamma_{\tau,N}(\tau_{N})}\right)\otimes I_{m},$
$B_{2,3}(\rho)$ $\coloneqq\text{diag}\left(\sqrt{1-\gamma_{\rho,1}(\rho_{1})},\sqrt{1-\gamma_{\rho,2}(\rho_{2})},...,\sqrt{1-\gamma_{\rho,{M^{\ast}}}(\rho_{M^{\ast}})}\right)\otimes I_{mN},$
$B_{2}(\theta)$ $\coloneqq\text{diag}\left(\sqrt{\epsilon}I_{n(N-1)},B_{2,2}(\tau),B_{2,3}(\rho)\right).$
Note that $B_{1}(\theta)v=0_{\varrho}$ and $B_{2}(\theta)v=0_{\varrho}$ cannot hold simultaneously as $B_{1}(\theta)+B_{2}(\theta)$ possesses the trivial null space. Consequently, $\mathsf{M}(0_{N},0_{M^{\ast}})<\mathbf{0}$, $\mathsf{M}(T_{2},T_{4})<\mathbf{0}$, and (52) imply $v^{\top}\mathsf{M}(\tau,\rho)v<0$ for all $v\neq 0_{\varrho}$ and $(\tau,\rho)\in\mathcal{T}\times\mathcal{R}$. ∎
Remark 1.
Corollary 1 provides a convenient method for validating the sufficient condition of Theorem 1. In particular, Corollary 1 only requires the checking of two matrix inequalities, namely, $\mathsf{M}(0_{N},0_{M^{\ast}})<\mathbf{0}$ and $\mathsf{M}(T_{2},T_{4})<\mathbf{0}$, to ensure $\mathsf{M}(\tau,\rho)<\mathbf{0}$ holds for all $(\tau,\rho)\in\mathcal{T}\times\mathcal{R}$. Nevertheless, Corollary 1 is not constructive in the sense that it does not provide a method for selecting $\sigma$, $T_{1}^{p}$, $T_{2}^{p}$, $T_{3}^{r}$, $T_{4}^{r}$, $K_{u}$, $K_{\eta,p}$, and $K_{\zeta,pr}$ for all $p\in\mathcal{V}$ and $r\in[M^{\ast}]$ so that $\mathsf{M}(0_{N},0_{M^{\ast}})<\mathbf{0}$ and $\mathsf{M}(T_{2},T_{4})<\mathbf{0}$ hold. Furthermore, Corollary 1 does not examine how the structure of the clustered network affects the set of matrices used to define $P(\tau,\rho)$ that satisfy $\mathsf{M}(0_{N},0_{M^{\ast}})<\mathbf{0}$ and $\mathsf{M}(T_{2},T_{4})<\mathbf{0}$. Since these topics will require an in-depth analysis, we reserve them for future work.
In Section 7, we will leverage Theorem 1 and Corollary 1 in an application and demonstrate that a set of parameters satisfying $\mathsf{M}(0_{N},0_{M^{\ast}})<\mathbf{0}$ and $\mathsf{M}(T_{2},T_{4})<\mathbf{0}$ exist.
6 Robustness
In this section, we briefly discuss the robustness properties of the set $\mathcal{A}$ for the hybrid system $\mathcal{H}$. Since $\mathcal{H}$ is nominally well-posed and the set $\mathcal{A}$ is compact and GES for $\mathcal{H}$, standard arguments in [Goebel.Sanfelice.ea2012, Chapter 7] and [Phillips.Sanfelice2019, Section 3.5.1] can be used to show that $\mathcal{A}$ is robust to vanishing perturbations for $\mathcal{H}$. For the remainder of this section, we examine the robustness of $\mathcal{A}$ to non-vanishing measurement noise, as defined in Definition 2, injected via a modified jump map of $\mathcal{H}$.